You are an investment banker trying to value Dell, a private company. You have forecasted Dell’s free cash flows, but need to compute its WACC in order to value the firm. Unfortunately Dell is private and so it does not have stock data, so you cannot use CAPM to find its cost of equity. You know the following: Dell has debt of $200 at a cost of 5%; further, Dell recently raised money from equity investors, valuing the equity at $1,000. Further, HP is in the same exact business as Dell, but it is public so you can see its cost of equity. HP is financed with a constant debt-to-equity ratio of 1/10, has a cost of debt of 3%, a cost of equity of 22%, and a tax rate of 30%.

1.Find the unlevered cost of capital for Dell and HP. Assume that HP’s debt-to-equity ratio will stay constant forever.

2.Using the unlevered cost of capital for Dell above, find the cost of levered equity for Dell (assuming that Dell’s capital structure D/E will remain fixed).

3.Using your above 2 answers, find Dell’s WACC assuming it has the same tax rate as HP.
Answer #1
1] We have
Levered cost of capital [rsl] = Unlevered cost of capital [rsu]+(unlevered cost of capital [rsu]-Cost of deb [rd]t)*D/E
Substituting values, we have
22 = rsu + (rsu-3)*0.1
Solving for rsu,
22 = rsu + 0.1*rsu-0.3
1.1*rsu = 22.3
rsu [unlevered cost of capital] = 22.3/1.1 = 20.27%
2] Cost of levered equity = 20.3+(20.3-5)*200/1000 = 23.33%
3] WACC = 23.33%*1000/1200+5%*(1-30%)*200/1200 = 20.03%